### v1

# Entanglement Microscopy: Tomography and Entanglement Measures via Quantum Monte Carlo

Abstract
We develop a protocol, dubbed entanglement microscopy, to obtain the full reduced density matrix associated with subregions in quantum Monte Carlo simulations for bosonic and fermionic many-body systems. Our microscopy allows to perform quantum state tomography, and thus gives access to true entanglement measures, such as the logarithmic negativity (LN). We exemplify our method by studying the phase diagram near quantum critical points (QCP) in 2 spatial dimensions: the transverse field Ising model and a Gross-Neveu-Yukawa transition of Dirac fermions. Our main results are: i) the Ising QCP exhibits short-range entanglement with a finite sudden death of the LN both in space and temperature; ii) the Gross-Neveu QCP has a power-law decaying fermionic LN consistent with conformal field theory (CFT) exponents; iii) going beyond bipartite entanglement, we find no detectable 3-party entanglement in a large parameter window near the Ising QCP in 2d, in contrast to 1d. We also analytically obtain the large-temperature power law scaling of the fermionic LN for general interacting systems. Our approach allows one to perform quantum state tomography locally in a way that is analogous to atomic-scale imaging with a scanning tunneling microscope. Controlled entanglement microscopy opens a new window into quantum matter, with countless systems waiting to be explored.
I Introduction
Quantum entanglement can reveal the fundamental organizing principles of quantum matter, ranging from non-Fermi liquids in high-temperature superconductors to black holes [1, 2]. However, obtaining the entanglement structure of realistic systems is a challenging task. When dealing with physically relevant situations that include finite temperature or the study of microscopic subregions, one has to work with mixed states (density matrices), and simple quantities like the von Neumann or Rényi entropies cannot be used as measures of entanglement. In the simplest case of bipartite entanglement, one can for example exploit the partial transpose to construct the (logarithmic) negativity [3, 4, 5], which is a true entanglement measure not polluted by classical correlations. However, it is challenging to compute in many-body systems for subregions with many degrees of freedom (including quantum field theories) where one can only obtain related quantities via the replica trick [6, 7]. These Rényi analogs are not entanglement measures and thus cannot be used to obtain the entanglement structure. A striking example is in boson/spin systems poised near a quantum critical point (QCP), where the logarithmic negativity (LN) decays faster than any power, as shown for CFTs in 1d [7] and higher $d$ [8], while the Rényi moments decay algrebraically. These are special cases of a general phenomenon regarding the “fate of entanglement” [9] where any multi-partite form of entanglement (including but more general than bipartite) entirely disappear as a system heats up, evolves in time, or as its parts grow separated.
To reveal the true microscopic structure of entanglement, in a way analogous to what the scanning tunneling microscope does for imaging atomic-scale densities, we use quantum Monte Carlo simulations to perform full tomography on the state of microscopic subregions. More precisely, we use a path-integral QMC approach that allows us to directly sample the reduced density matrix (RDM) of various subregions, as shown in Fig. 1. Whereas a scanning tunneling microscope usually works with one tip, here we exploit numerous “tips” to capture $m$-party entanglement. In the present paper, we illustrate such entanglement microscopy for $m=2,3$.
Specifically, we use entanglement microscopy to scrutinize systems near generic 2d quantum phase transitions: first the celebrated QCP in the transverse field Ising model (TFIM) on the square lattice, and second a Gross-Neveu-Yukawa QCP of interacting Dirac fermions. Regarding bipartite entanglement, our tomography allows the computation of the LN in the entire phase diagram including at finite temperature with the main findings: i) the Ising model exhibits short-range entanglement with finite sudden death as a function of space and temperature; ii) the fermionic variant of the LN, which takes into account the parity superselection rule, shows power-law decay at the Dirac QCP for large separations with exponents consistent with the conformal field theory (CFT) prediction [8]. iii) Going beyond bipartite entanglement, we do not detect any genuine 3-party entanglement in a large window near the 2d Ising QCP, but find its presence in the 1d model both in QMC and with the exact solution via fermionization. Overall, the 2d Ising model shows less bipartite and tripartite entanglement compared to 1d [10, 11, 12, 13];
Our entanglement microscopy scheme offers advantages in terms of immense versatility and computational efficiency. By focusing on microscopic or so-called “skeletal” subregions [14], entanglement microscopy not only gives the full RDM and true entanglement measures, but it does so without the need of replicating the entire system [15], which leads to increases in computational complexity, numerical instabilities even with advanced algorithms [16, 17, 18, 19, 20, 21, 22, 23], and stringent requirement of data quality during the analytic continuation procedures [24].
Our paper is organized as follow: in Section II we introduce the entanglement microscopy method in QMC. In Section III we discuss how true entanglement measures can be extracted between 2 and 3 parties. In Section IV we present the QMC results for the Ising and Dirac models. We also present an analytical result for the large-temperature scaling of the fermionic LN in general many-body fermion systems. Finally, we give a conclusion in Section V, where we outline the countless future applications of entanglement microscopy.
Figure 1: Entanglement microscopy: sampling the reduced density matrix in QMC. Panel (a) shows the partition function in the space ($r$) - time ($\tau$) manifold appearing in usual QMC simulations, with the periodic structure of imaginary time. Panel (b) shows how the RDM is sampled in such QMC simulations, where one imposes open boundary conditions in time for the skeletal sites in subregion $A$. From the QMC histogram of configurations of these sites, in this case, there will be $4\times 4$ configurations for the 2 spins in $A$, the corresponding matrix elements in the RDM are recorded. The lower panel in (b) demonstrates the spatial partition of $A$ and $B$ in such a setting.
II Entanglement microscopy
The entanglement microscopy protocol is illustrated in Fig. 1, in the realization of the path-integral QMC for quantum spin systems. In such stochastic series expansion (SSE) QMC simulations [25], the RDM of the spins in subregion $A$, $\rho_{A}$, is evaluated as
$\begin{split}\rho_{A}&=\frac{1}{Z}\operatorname{Tr}_{B}\{e^{-\beta H}\}=\frac{1}{Z}\sum_{\xi}\bra{\xi}e^{-\beta H}\ket{\xi}\\ &=\frac{1}{Z}\sum_{\{b_{1},b_{2}\cdots b_{M}\}}\frac{\beta^{n}(M-n)!}{M!}\sum_{\xi}\bra{\xi}H_{b_{M}}\cdots H_{b_{2}}H_{b_{1}}\ket{\xi},\end{split}$ (1)
where {|$\{\ket{\xi}\}$ forms a complete basis of the environment $B$. The sampling of the RDM is done in an enlarged ensemble where one does not enforce the left state and right state to be the same in region $A$, i.e., we discard the periodic boundary condition in time. As shown in Fig. 1 (b), by keeping the periodic boundary condition in region $B$ and open boundary in region $A$, in each Monte Carlo step, one records the spin configuration on each site in region $A$, and the frequency of each element appearing is proportional to its corresponding element in the RDM.
For interacting fermions, one employs the determinant quantum Monte Carlo (DQMC) scheme [26]. In DQMC, the sampling of the RDM is attributed to the fact that fermions are considered free in each auxiliary field configuration, thus the interacting RDM is the weighted average of free RDMs [27]. For the case where region $A$ only contains two sites, the interacting RDM in the occupation-number basis can be constructed with the density correlation function, and fermion Green’s function [8].
We note that DQMC calculation of Rényi negativity with bipartition at finite temperature of interacting fermions, via the replica trick, has recently been worked out [28]. However, since the RDM therein is not explicitly computed, only integer moments of LN can be obtained.
Figure 2: Logarithmic negativity for adjacent sites. Color code plots of (a) $\mathcal{E}^{b}$ in 2d TFIM simulated in a $40\times 40$ square lattice with $\beta$ up to 10, and (b) $\mathcal{E}^{f}$ in 2d fermion $t-V$ model simulated in a $6\times 6$ square lattice with $\beta$ up to 15. In both panels, the red curves represent the finite-temperature phase boundaries which end at a zero temperature QCP located at $h_{c}=3.044$ in 2d TFIM [29] and $V_{c}=1.27$ in $t-V$ model. The purple dashed lines show the maxima of $\mathcal{E}^{b/f}$ for each temperature. In both models, the maxima have converged with system size, as is shown in the SM [30]. The bottom panel illustrates the schematic phase diagrams.
III Entanglement measures & Multipartite entanglement criteria
A strong advantage of our entanglement microscopy is the versatility of the full RDM: we can use it to compute not only the expectation value of any observable but also any entanglement measure involving the degrees of freedom in $A$. An example is the logarithmic negativity that can be used to shed light on bipartite entanglement, and its evolution in the entire phase diagram [14, 8]. More interestingly, we can also evaluate various multipartite entanglement criteria, and these reveal more collective forms of entanglement.
III.1 Bipartite entanglement
The logarithmic negativity is calculated with RDM $\rho_{A}$ where region $A$ is subsequently divided into two part, $A_{1}$ and $A_{2}$, and one takes the partial transpose in region $A_{1}$. It is defined as,
$\mathcal{E}^{b/f}=\ln||\rho_{A}^{T^{b/f}_{1}}||_{1}$ (2)
[…]
III.2 Tripartite entanglement
In the above, we have discussed bipartite entanglement between two sites. Entanglement microscopy allows us to delve into the much richer realm of multi-party entanglement. We illustrate this by studying tripartite entanglement between 3 sites in the 2d quantum Ising model. In particular, we shall search for genuine tripartite entanglement (GTE), which cannot be described in terms of 2-party entanglement between any two subgroups of spins. A state with GTE cannot be written as a mixture of bi-separable physical density matrices, $\rho\neq\sum_{k}p_{k}\rho_{A}^{k}\rho_{BC}^{k}+q_{k}\rho_{AB}^{k}\rho_{C}^{k}+r_{k}\rho_{AC}^{k}\rho_{B}^{k}$, where $p_{k},q_{k},r_{k}$ are non-negative. A simple example of GTE is given by a pure cat state (also known as GHZ state) of 3 spins, $a|000\rangle+b|111\rangle$. Criteria to detect GTE are more involved than those of the simple bipartite measures such as the LN. However, some criteria exist that can detect GTE by establishing that a state is not bi-separable. Here, we shall focus on two such separability criteria, that we call $W_{1}$ and $W_{2}$ [35]. Let us describe $W_{1}$ ($W_{2}$ is analogously defined, see Supplemental Material (SM) [30]):
$W_{1}\!=\!\max_{\,\rm LU}\left(\left|\rho_{1,8}\right|-\sqrt{\rho_{2,2}\rho_{7,7}}-\sqrt{\rho_{3,3}\rho_{6,6}}-\sqrt{\rho_{4,4}\rho_{5,5}}\right)$ (3)
where $\rho_{i,j}$ are the matrix element of the 3-site RDM. In order to get a meaningful answer, one has to remove the basis dependence by maximizing among all local unitary transformations acting on the RDM. Being local, such transformations leave the GTE structure invariant. If $W_{i}>0$, there is GTE, while no conclusion can be reached if $W_{i}\leq 0$. Although not sensitive to all forms of GTE, the $W_{i}$ criteria do tend to perform well in a variety of states, especially when used together [35].
Figure 3: Spatial and temporal dependence of the logarithmic negativity. (a) The decay of the two-site $\mathcal{E}^{b}$ and $\mathcal{E}^{b}_{3}$ as a function of separation $r$ at the QCP $h_{c}=3.04$. The simulation is done with $L=\beta=40$. $\mathcal{E}^{b}$ exhibits a sudden death within one lattice constant, and the $\mathcal{E}^{b}_{3}$ exhibits a power-law decay with the fit shown as a purple dashed line; the exponent $3.9(2)$ is consistent with the expected Ising CFT value $4\Delta_{\sigma}+2\approx 4.073$ [8, 36]. (b) The decay of $\mathcal{E}^{f}$ versus separation $r$ in the $t-V$ model with various system sizes, at the QCP $V_{c}=1.27$. The simulation is done with $L=\beta$ going from 6 to 14. The blue dashed line shows the algebraic scaling with the power predicted by CFT analysis [8, 37], with the expected GNY scaling dimension $4\Delta_{\psi}\approx 4.174$. Panel (c) and (d) demonstrate the temperature dependence of $\mathcal{E}^{b}$ for the $L=40$ TFIM, and $\mathcal{E}^{f}$ for the $L=6$ $t-V$ model at various values of $h$ and $V$. In (c) the $\mathcal{E}^{b}$ always exhibit a finite sudden death temperature and in (d) $\mathcal{E}^{f}$ decays as $1/T^{2}$ for adjacent sites $r=1$, while $1/T^{6}$ for sites separated by $r=3$ bonds.
IV Results
IV.1 Models
We perform entanglement microscopy on the 1d and 2d square lattice TFIM, as well as the spinless $\pi$-flux fermion $t-V$ model on a square lattice. The Hamiltonian of the TFIM reads
$H=-J\sum_{\langle i,j\rangle}\sigma^{z}_{i}\sigma^{z}_{j}-h\sum_{i}\sigma^{x}_{i},$ (4)
where we have set the ferromagnetic exchange coupling to $J=1$; the QCP is at $h=1$ for 1d, and $h=3.044$ for 2d [29]. The $T$-$h$ phase diagram is shown in panel (a) of Fig. 2. As will be shown below, besides the QMC simulation for RDMs, we have also performed exact diagonalization (ED) in 2d for small clusters, to obtain both bipartite and tripartite entanglement measures.
For the spinless fermion $t-V$ model, the Hamiltonian reads,
$H=-t\sum_{\langle i,j\rangle}(e^{i\theta_{i,j}}c^{\dagger}_{i}c_{j}+{\rm h.c.})+V\sum_{\langle i,j\rangle}(n_{i}-\tfrac{1}{2})(n_{j}-\tfrac{1}{2}),$ (5)
it consists of nearest neighbor hopping term with $\pi$-flux and nearest neighbor repulsion. We set $t=1$ as the energy unit. The $\pi$-flux hopping term on the square lattice ($\theta_{i,j}=\frac{\pi}{2}$ on every other column of vertical bonds and 0 otherwise) leads to a dispersion with two Dirac cones at the $X=(\pi,\pm\frac{\pi}{2})$ points in the Brillouin zone (BZ) taking two sites horizontally as a unit cell. At half filling, a positive $V$ drives the system from a Dirac semi-metal (DSM) to a $\mathbb{Z}_{2}$-symmetry spontaneously broken charge density wave (CDW) state, where one sublattice acquires a greater density compared to the other. The QCP is described by the “chiral Ising” Gross-Nevey-Yukawa (GNY) theory with global symmetry $O(N_{f})^{2}\rtimes\mathbb{Z}_{2}$ and in our case $N_{f}=2$ two-component Dirac fermions. The position of the QCP, the critical exponents of the Gross-Neveu transition, have been systematically worked out using QMC, field-theoretical expansions, and conformal bootstrap, for both $N_{f}=2$ [38, 37, 39, 40, 41, 42], and other values of $N_{f}$ [43, 37, 39, 40, 44, 41, 42, 45, 46, 47, 36, 48, 49, 50]. The schematic $T-V$ phase diagram is shown in lower part of the panel (b) in Fig. 2.
By simulating this model with $L=\beta=6,8,10,12,14$, we can probe the chiral GNY-Ising QCP. Here the linear size $L$ gives the number of unit cells. We located the QCP at $V_{c}=1.27$ via finite-size scaling and using the critical exponents from conformal bootstrap [37]. The detailed QMC finite-size data analysis and collapse are shown in the SM [30].
IV.2 Bipartite entanglement
We now present the results for $\mathcal{E}^{b}$ of the TFIM and $\mathcal{E}^{f}$ of the $t-V$ model.
IV.2.1 Ising model
Fig. 2 (a) shows the nearest neighbor $\mathcal{E}^{b}$ at finite temperature for the square lattice TFIM of linear system size $L=40$ with inverse temperature $\beta$ up to 10. The maxima of $\mathcal{E}^{b}$ at each temperature, indicated by the purple dashed line, go from the disordered state at high temperature to $h=3.1$ at the lowest temperature $T=1/10$, which is close to but distinct from the QCP, $h_{c}=3.044$. At all $h$ values, $\mathcal{E}^{b}$ has a sudden-death temperature, beyond which $\mathcal{E}^{b}$ is identically zero corresponding to the dark blue region in the phase diagram; it is also shown in panel (c) of Fig. 3. The existence of a thermal sudden death follows from the general fact that the infinite temperature state lies well within the continent of separable states [9], and so entanglement must disappear at a finite temperature corresponding to the point at which the system enters the continent.
Let us now examine the fate of entanglement as the 2 sites become progressively more separated. Fig. 3 (a) shows the $\mathcal{E}^{b}$ between two sites separated with bond-distance $r=|i-j|$. $\mathcal{E}^{b}$ drops to 0, within machine precision, even for the next-nearest neighbor separation, $r=\sqrt{2}$. Strikingly, the 2d TFIM possesses the shortest possible range of entanglement for two spins, as it only exists for nearest neighbor, even shorter than the 1d case where the sudden-death distance was found to be $r_{\textrm{sd}}=3$ [51, 11, 13, 30]. Such a sudden death separation again follows since the state at large separations lies within the separable continent [9].
We also computed the third order moment $\mathcal{E}^{b}_{3}$ at the Ising QCP. The moments are defined [6, 7] as $\mathcal{E}^{b}_{n}=\ln\frac{\operatorname{Tr}\rho_{A}^{n}}{\operatorname{Tr}[(\rho_{A}^{T_{1}})^{n}]},$ which can be easily computed with the RDM obtained via microscopy, or one can sample it by using standard QMC and exploiting the replica trick [52]. The orange line in the inset of Fig. 3 (a) shows the power-law decay of $\mathcal{E}^{b}_{3}$ with fitted power 3.9(2), which agrees with the CFT prediction $4\Delta_{\sigma}+2\approx 4.073$ [8, 36]. The power-law decay of $\mathcal{E}^{b}_{3}$ is in sharp contrast with the rapid sudden death of $\mathcal{E}^{b}$, and arises since the former is not a proper entanglement measure, and is thus polluted by non-entangling correlations.
IV.2.2 Fermions
Fig. 2 (b) shows the analogous results for $\mathcal{E}^{f}$ for the square lattice $t-V$ model of linear system size $L=6$ with $\beta$ up to 15. As denoted by the purple dashed line, the maxima starts in the ordered CDW phase at high temperature and converges to $V\approx 1$ at low temperature, which is in the DSM phase. It roughly follows the finite-temperature phase boundary, obtained from finite-size scaling [30] and indicated by the red line, but clearly differs from it at most temperatures. At high $T$, the nearest neighbor $\mathcal{E}^{f}$ does not suffer a sudden death. Instead, it decays algebraically as shown in Fig. 2 (b) and Fig. 3 (d).
We can actually obtain the scaling analytically in the large $T$ limit for general models and separations. In order to determine $\mathcal{E}^{f}$ in a thermal state $\rho=\exp(-\beta H)/\mathcal{Z}$, we need to find how the fermion Green’s function $g_{ij}=\operatorname{Tr}(c^{\dagger}_{i}c_{j}\rho)$ decays to zero. We will then have $\mathcal{E}^{f}=\alpha|g_{ij}|^{2}+\cdots$, where the coefficient $\alpha$ depends on the filling, and the ellipsis denote subleading terms. By expanding $\exp(-\beta H)=\sum_{n}(\beta H)^{n}/n!$ we see that the first non-zero contribution to the Green’s function will come from $\operatorname{Tr}(H^{n_{r}}c^{\dagger}_{i}c_{j})$ such that the hopping term of $H^{n_{r}}$ connects sites $i$ and $j$. The Green’s function thus decays as $1/T^{n_{r}}$, and we have
$\mathcal{E}^{f}=A\,T^{-2n_{r}}+\cdots$ (6)
To illustrate the result, let us consider a general Hamiltonian that includes nearest neighbor hopping terms, an example being our $t$-$V$ model. If the sites are adjacent, $n_{r}=1$ and we get $1/T^{2}$. This scaling is in agreement with the result for adjacent regions in general Gaussian (free) fermionic states [32], but we emphasize that it holds in interacting systems. If the sites are not adjacent, we will get $1/T^{2n_{r}}$ scaling with $n_{r}\geq 1$. For instance, in the $t$-$V$ model for third nearest neighbor sites, $n_{r}=3$, leading to $1/T^{6}$ scaling, as shown in Fig. 3 (d). Further information about the fits are given in the SM [30]. Finally, the analysis above can be generalized to subregions with more sites: we infer that the exponent is the minimal one for all combinations of sites $i\in A_{1}$ and $j\in A_{2}$. Interestingly, a smaller exponent $n=3/2$ was observed in topological integer quantum Hall states for adjacent regions, [53] which does not contradict our general analysis since quantum Hall states live in the continuum.
Finally, Fig. 3 (b) shows the scaling of $\mathcal{E}^{f}$ as we increase the separation between the two sites at the GNY QCP. In contrast to $\mathcal{E}^{b}$ for the Ising QCP, the fermionic LN shows long-range entanglement, manifested by its power-law decay with $r$. The dashed line shows the CFT prediction [8] with power $4\Delta_{\psi}\approx 4.174$, where the fermion scaling dimension was obtained via conformal bootstrap [37]. The algebraic tail of the fermionic LN follows from the general structure of the space of separable states in fermion systems owing to the parity superselection rule [9].
Figure 4: Searching for genuine tripartite entanglement in the 1d/2d quantum Ising models. In 1d, the criterion $W_{1}\geq 0$ detects GTE for 3 consecutive spins at all the values of $h$. In 2d for 3 spins in a square plaquette, $W_{1}\leq 0$, both in the small size exact diagonalization (ED) and large-scale QMC results, signifying the absence of the GTE in a large region near the QCP. Similar behavior for $W_{2}$ is shown in the SM [30]. All QMC calculations are done with $\beta=L$.
IV.3 Tripartite entanglement
In 1d, GTE of three consecutive spins was found in the 1d anisotropic XY model, which includes the Ising case as a special limit [12], using a quantum $k$-separability criterion [54]. GTE has also been studied in a cluster-Ising model that gives rises to RDMs with special fine-tuned properties [55], while multi-partite entanglement of the entire system, not subregions as considered here, was studied for small system sizes in the TFIM in dimensions 1 to 3 [56]. Quantum microscopy of the GTE beyond 1d has not been previously performed. Fig. 4 shows the GTE criterion $W_{1}$ for the three closest sites as a function of the transverse field $h$ for both the 1d and 2d TFIMs. The 1d data, obtained from the exact RDM via a Jordan-Wigner transformation [57], shows that GTE exists for all finite values $h$, and becomes maximal near the QCP. In fact, $dW_{1}/dh$ diverges at QCP, exactly as was found for a related criterion $dI_{2}/dh$ in [12]. In sharp contrast, in the 2d case, the large-scale QMC data has $W_{1}\leq 0$ for all values $h$, indicating that no GTE is detected. We have also evaluated another complementary criterion $W_{2}$, and reached the same conclusion in the range $20<h/h_{c}<$ 2 (see SM [30]). Furthermore, we have checked that in both 1d and 2d no GTE is detected if the sites are non-adjacent. Extrapolating to higher dimensions, we expect that GTE decreases with increasing $d$. In particular, for the 3d Ising model, we expect that $W_{i}<0$ in a large window near the QCP. It will be interesting to test our prediction using the entanglement microscopy methods that we have developped.
Interestingly, in 2d, using ED on clusters up to 25 sites, we found that a small $W_{2}\sim 2\times 10^{-4}$ appears at large values of the transverse field $h\gtrsim 8$, and our finite-size scaling analysis suggest that it will survive in the thermodynamic limit (see SM [30]).
V Discussion
In this work, we have put forward a framework – entanglement microscopy – to perform full tomography within QMC, giving access to the true entanglement hidden in quantum many-body systems. To demonstrate the power of our scheme, we study the phase diagram of two representative QCPs in 2 spatial dimensions: the transverse field Ising model and the GNY chiral-Ising transition of Dirac fermions. We find i) the Ising QCP exhibits short-range entanglement with a finite sudden death as a function of separation and temperature; ii) the fermionic QCP exhibits power-law decay of the fermionic LN at large separations between the subregions; iii) there is no detectable 3-party entanglement in a large window near the 2d Ising QCP, while it is present in the 1d counterpart. Our results demonstrate that controlled entanglement microscopy is now possible on large-scale quantum many-body systems in any dimension. A new window into quantum many-body entanglement has opened, with countless systems waiting to be explored.
In the future, it will be interesting to explore the multipartite entanglement at various exotic QCPs, including transitions into a non-Fermi-liquid [58, 59], deconfined QCPs for both fermion and spin models [60, 61, 62, 63, 64, 65, 66]. In such explorations, the recent realisation that the very structure of the space of physical states determines the fate of entanglement as a system evolves will be useful to guide entanglement microscopy in realistic many-body Hamiltonians [9].
Note added.— During the completion of this manuscript, we became aware of a work that shared a similar idea on sampling RDM in spin models [67].
Acknowledgments.—We acknowledge discussions with Gilles Parez. T.T.W., M.H.S. and Z.Y.M. acknowledge the support from the Research Grants Council (RGC) of Hong Kong Special Administrative Region (SAR) of China (Project Nos. 17301721, AoE/P701/20, 17309822, C7037-22GF, 17302223), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A HKU703/22) and the HKU Seed Funding for Strategic Interdisciplinary Research. W.W.-K. and L.L. are supported by a grant from the Fondation Courtois, a Chair of the Institut Courtois, a Discovery Grant from NSERC, and a Canada Research Chair. We thank HPC2021 system under the Information Technology Services and the Blackbody HPC system at the Department of Physics, University of Hong Kong, as well as the Beijng PARATERA Tech CO.,Ltd. (URL: https://cloud.paratera.com) for providing HPC resources that have contributed to the research results reported within this paper.
Supplemental Material for
"Entanglement Microscopy: Tomography and Entanglement Measures via Quantum Monte Carlo"
In this supplementary material, we will provide some additional data on 1d and 2d TFIM in Sec. I, and 2d fermion $t-V$ model in Sec. II.
In Sec. I, we will show the 1d result on the LN $\mathcal{E}^{b}$ and systematic comparison with ED in addition to the 2d result, which has been covered in the main text. We have also done a finite size convergence test in several parameter settings.
In Sec. II, we will show some observables we collected in the $t-V$ model with $L=\beta$. Finite-size scaling, single-particle gap scaling and Green’s function decay are done to justify the location of QCP. LN in this setting is also computed and shown. Then, we will show the finite-temperature data including a similar convergence test and finite size scaling to locate the finite-temperature CPs. Lastly, a color code plot of $\mathcal{E}^{b}$ is presented.
I 1d & 2d Transverse Field Ising Model
Figure S1: 2-site entanglement negativity with separation $r$ of 1d and 2d transverse field Ising models. Log negativity $\mathcal{E}^{b}$ (a), and its third order moment $\mathcal{E}^{b}_{3}$ (b) in 1d Ising model with system size and inverse temperature $L=\beta=128$, and $h=1$. Log negativity $\mathcal{E}^{b}$ (c), and its third order moment $\mathcal{E}^{b}_{3}$ (d) in 2d Ising model with system size and inverse temperature $L=\beta=40$, and $h=3.044$.
Fig. S1 (a) and (b) show the decay of $\mathcal{E}^{b}$ and $\mathcal{E}^{b}_{3}$ against the separation distance between two sites for the (1+1)d Ising model on the QCP. The logarithmic negativity (LN) $\mathcal{E}^{b}$ drops abruptly to $10^{-16}$ (machine precision) with sudden-death distance $r_{\textrm{sd}}=3$ [13], while its third moment decay algebraically with fitted power $2.49(3)$, which agrees with both exact calculation using Jordan Wigner Transformation and CFT prediction $2.52+4\Delta_{\sigma}=$ 2.5 [8].
Similar to the case in 1d TFIM, as shown in Fig. S1 (c), the LN $\mathcal{E}^{b}$ drop to the magnitude of machine error with sudden-death distance $r_{\textrm{sd}}=\sqrt{2}$, i.e., only nearest neighbor entangle exists. Its third order moment $\mathcal{E}^{b}_{3}$ decay algebraically with fitted power $3.9(2)$, where CFT prediction is $4\Delta_{\sigma}+2\approx 4.073$ [8, 36].
Figure S2: Nearest neighbor LN $\mathcal{E}^{b}$ at various system sizes. (a) $\mathcal{E}^{b}$ changes against the transverse field $h$ at different temperature and sizes. For each temperature, $\mathcal{E}^{b}$ is converged against system size. The triangles indicate the maxima of $\mathcal{E}^{b}$ along the $h$ axis (the purple dots in Fig. 2 in the main text). (b) $\mathcal{E}^{b}$ changes against the inverse temperature $\beta$ at different transverse field $h$ and sizes. For each $h$, $\mathcal{E}^{b}$ is also converged against system size.
Fig. S2 shows the convergence test of the nearest neighbor LN $\mathcal{E}^{b}$ against system size $L$. In all cases, the results from these three sizes show good convergence in terms of size. This strongly suggests that the $L=40$ data presented in main text has negligible finite-size effect.
I.1 Comparison with Exact Diagonalisation
Figure S3: Entanglement measures from ED. (a) The decay of the two-site $\mathcal{E}^{b}$ against the separation distance $r$, with $L_{x},L_{y}=5,5$, $T=0$ and $h=3.04$ (b) Two-site $\mathcal{E}^{b}$ against transverse field h for two sites at distance $r=1$ and $\sqrt{2}$, with $L_{x},L_{y}=4,5,T=0$. (c) The decay of two-site $\mathcal{E}^{b}$ with temperature at distance $r=1$ for $h=2,3,6$, with $L_{x},L_{y}=4,4$, $T=0$. (d) Two GTE criteria against transverse field $h$, with $L_{x},L_{y}=4,5$, $T=0$. The three sites form the smallest right-triangle. The inset show $W_{2}$ going negative for large $h$.
Figure S4: GTE at large $h$ scaling with size. GTE criterion $W_{2}$ has a small positive value of the order $10^{-4}$ at large $h$. We show how this value scales with the size of the system, where $N=L_{x}L_{y}$ is the number of spins. For $N=5\times 5$, $W_{2}$ appears to have converged.
For small system sizes, we can obtain RDM through exact diagonalisation (ED) and compare with the QMC results. In this section, we compute different entanglement measures using ED. Apart from some finite size effect, these results agree with that of QMC.
I.1.1 $T=0$
At $T=0$, RDM is constructed from the ground state $\psi_{0}$ only: $\rho_{A}=\operatorname{Tr}_{B}\{|\psi_{0}\rangle\langle\psi_{0}|\}$. In Fig. S3 (a),(b),(d), we show how the two-site $\mathcal{E}^{b}$ change with distance $r$ and field $h$, and how the GTE critera $W_{1}$, $W_{2}$ change with field $h$.
Fig. S3 (a) shows the decay of the two-site $\mathcal{E}^{b}$ against the separation distance $r$, which can be compared with the $L=40$ QMC result in Fig. S3 (a). We see that $\mathcal{E}^{b}(1)\approx 0.11$ in both cases, but the ED $\mathcal{E}^{b}(r)$ drop to zero (within machine precision) at distance $r=2$, while the QMC result drops to zero already at distance $r=\sqrt{2}$. This is likely a finite size effect, as with ED we are limited to $L=5$.
Fig. S3 (b) shows the two-site $\mathcal{E}^{b}$ as a function of the transverse field $h$. At both distance $r=1$ and $r=\sqrt{2}$, $\mathcal{E}^{b}$ peak at around $h=3$, close to the QCP. At large $h$, $\mathcal{E}^{b}(1)\sim 1/h^{0.9}$ and $\mathcal{E}^{b}(\sqrt{2})\sim 1/h^{4.6}$. For small $h$, $\mathcal{E}^{b}(\sqrt{2})$ has a sudden death at $h\sim 0.5$.
Fig. S3 (d) shows two GTE criteria $W_{1},W_{2}$ against $h$. The two criteria are defined in Ref. [35] as
$W_{1}[\rho]$ $=\max_{\,\rm LU}\left(\left|\rho_{1,8}\right|-\sqrt{\rho_{2,2}\rho_{7,7}}-\sqrt{\rho_{3,3}\rho_{6,6}}-\sqrt{\rho_{4,4}\rho_{5,5}}\right)$
$W_{2}[\rho]$ $=\max_{\,\rm LU}\left(\left|\rho_{2,3}\right|+\left|\rho_{2,5}\right|+\left|\rho_{3,5}\right|-\frac{1}{2}\left(\rho_{2,2}-\rho_{3,3}+\rho_{5,5}\right)-\sqrt{\rho_{1,1}\rho_{4,4}}-\sqrt{\rho_{1,1}\rho_{6,6}}-\sqrt{\rho_{1,1}\rho_{7,7}}\right)$
where the two functions are maximized with respect to local unitary transformations $U=U_{1}\otimes U_{2}\otimes U_{3}$ to the reduced density matrix $\rho$. The two separability criteria are defined such that if $W_{i}>0$, there is genuine three-qubit entanglement. If $W_{i}\leq 0$, we cannot conclude anything. $W_{1}$ and $W_{2}$ are strong criteria for GHZ and W states, respectively. In Fig. S3 (d), we see that near $h\sim h_{c}$ both criteria are negative. For $h>7$, the inset shows that $W_{2}$ go positive for $h>7.6$, indicating a small GTE at large $h$. In Fig. S4, we see that $W_{2}$ seem to converge to a fixed value as we increase the size of the system at large $h$. This suggests that the small GTE at large $h$ should survive in the thermodynamic limit.
I.1.2 Finite T
At finite temperature $T$, RDM is constructed with the full energy spectrum: $\rho_{A}=\frac{1}{Z}\operatorname{Tr}_{B}\{\sum_{n}e^{-\beta E_{n}}|\psi_{n}\rangle\langle\psi_{n}|\}$. If $T$ is not high, the Boltzmann factor $e^{-\beta E_{n}}$ of high energy eigenstates are diminishing, so we can cut the summation to a finite number of lowest energy eigenstates. We found that for $30$ lowest eigenstates, $\mathcal{E}^{b}$ already converges to the values shown in Fig. S3 (c).
Fig. S3 (c) shows the decay of $\mathcal{E}^{b}(1)$ with temperature, which is very similar to the $L=40$ QMC result in Fig. S3 (c), in terms of the shape of the three curves and the magnitude. There is a small difference in the sudden death temperature, likely due to the finite size effect.
I.2 Example of RDM
Here we give an example result of 3-site RDM for the 2d TFIM with $L=\beta=24$ and $h=3$. The three sites are chosen the same way as we did in investigating tripartite entanglement, and form on the smallest right triangle with the first site on the right angle.
$\begin{pmatrix}0.22268(4)&0.13787(4)&0.13182(3)&0.10746(2)&0.13182(3)&0.10746(2)&0.11605(3)&0.12789(5)\\ 0.13787(4)&0.09911(4)&0.08899(3)&0.08646(2)&0.08899(3)&0.08646(2)&0.08746(4)&0.11605(3)\\ 0.13182(3)&0.08899(3)&0.08911(2)&0.08046(3)&0.08408(3)&0.07709(3)&0.08646(2)&0.10746(2)\\ 0.10746(2)&0.08646(2)&0.08046(3)&0.08911(2)&0.07709(3)&0.08408(3)&0.08899(3)&0.13182(3)\\ 0.13182(3)&0.08899(3)&0.08408(3)&0.07709(3)&0.08911(2)&0.08046(3)&0.08646(2)&0.10746(2)\\ 0.10746(2)&0.08646(2)&0.07709(3)&0.08408(3)&0.08046(3)&0.08911(2)&0.08899(3)&0.13182(3)\\ 0.11605(3)&0.08746(4)&0.08646(2)&0.08899(3)&0.08646(2)&0.08899(3)&0.09911(4)&0.13787(4)\\ 0.12789(5)&0.11605(3)&0.10746(2)&0.13182(3)&0.10746(2)&0.13182(3)&0.13787(4)&0.22268(4)\end{pmatrix}$
II Fermion $t-V$ model
II.1 The $N=4$ GNY quantum critical point
Figure S5: Observables of the $t-V$ model. (a) Order parameter $M^{2}$, its (b) finite-size scaling crossing points and (c) correlation ratio $R$, and (d) data collapse with the two observables. Simulation is done with inverse temperature set to equal system size $\beta=L$. All four panels share the same legend in (d).
In finding the QCP of $t-V$ model, we perform the finite-temperature DQMC in various system sizes and set $\beta$ scale linearly with system size, since the fermion dispersion is linear ($z=1$) at the QCP. Fig. S5 shows the order parameter $M^{2}=S(0,0)$, where $S(k_{x},k_{y})=\sum_{i,j,\alpha}e^{-i(k_{x},k_{y})\cdot{\mathbf{r}_{ij}}}(n_{i,\alpha}n_{j,\alpha}-n_{i,\alpha}n_{j,\bar{\alpha}})$, correlation ratio $R=1-\frac{S(2\pi/L,0)}{S(0,0)}$, the finite-size crossing of them and data collapse of $M^{2}$ using the critical exponents from Ref. [37].
Panel (b) shows crossing near $V=1.265$, drifting slightly towards larger $V$, which is a sign of small finite-size effect, while the crossing of correlation ratio, an RG invariant value by design, shows more severe drifting from $V=1.28$ towards smaller $V$. The data collapse of them, shown in panel (d), with the usual scaling form shows good collapse for different sizes with $V_{c}=1.27$.
Figure S6: Single-particle gap versus $V$. finite-size gap fitted using the decay of imaginary time Green’s function $G({\mathbf{k}},\tau)\propto e^{\Delta_{\mathbf{k}}}$ at Dirac points. (a) finite-size gap against $V$. (b) Scaling of finite-size gap against $1/L^{\nu}$. Dashed lines in panel (b) are straight lines fitted with the three largest system sizes.
With the above setting, we also calculated the imaginary time decay of the Green’s function $\tilde{G}({\mathbf{r}}_{i}-{\mathbf{r}}_{j},\tau)=\langle c_{i}(\tau)c^{\dagger}_{j}(0)\rangle$, and its Fourier component $\tilde{G}({\mathbf{k}},\tau)=\sum_{{\mathbf{r}}}e^{-i{\mathbf{r}}\cdot{\mathbf{k}}}\tilde{G}({\mathbf{r}},\tau)$. At each ${\mathbf{k}}$ point, the Green’s function decays exponentially as $\tilde{G}({\mathbf{k}},\tau)=\langle c_{k}(\tau)c^{\dagger}_{k}(0)\rangle\propto e^{\Delta_{\mathbf{k}}}$, which can therefore be used to calculate the single-particle gap at the Dirac points (${\mathbf{k}}=(\pi,0)$ and ${\mathbf{k}}=(0,\pi)$).
As shown in Fig. S6 (a), the gap remains 0 at small $V$ and opens at around $V=1.2$. Panel (b) shows the scaling of finite-size gap $\Delta$ against system size $L$ in the critical regime. At the QCP, the gap should close as $\Delta=1/L^{z\nu}$. The y-intercepts of the fitted straight dashed lines suggest that the gap opens near $V=1.27\sim 1.28$ in thermodynamic limit $L\rightarrow\infty$.
Figure S7: Imaginary time decay of the fermionic Green’s function $\tilde{G}^{f}(\tau)$. Dashed line is the power-law decay function with power $2+\eta_{\psi}$.
For the onsite imaginary time Green’s function $\tilde{G}^{f}(\tau)=\langle c_{i}(\tau)c^{\dagger}_{i}(0)\rangle$, it should decay algebraically at the QCP with power $2+\eta_{\psi}$. As shown in Fig. S7, the decays for all system sizes coincide with the power-law decay with scaling dimension $\eta_{\psi}$ from Ref. [37].
Figure S8: Nearest neighbor LN $\mathcal{E}^{f}$ against $V$. In all cases, inverse temperature $\beta$ is set to equal to the system size. The red star in panel is the ED result of free fermion model with system size and inverse temperature $L=\beta=1000$. Panel (b) is a zoom-in plot of the square region in (a). Panels (a) and (b) share the same legend in (a)
Fig. S8 shows the result of nearest neighbor $\mathcal{E}^{f}$ against repulsion strength $V$. For $V>V_{c}$, the system is in the gapped CDW phase, which explains the good convergence even in higher temperature. For $V<V_{c}$, the system is in the gapless DSM phase, and $\mathcal{E}^{f}$ is drifting up as $L$ and $\beta$ increase. The $L=\beta=12$ curve extrapolates to the ED result in $L=\beta=1000$, indicated by the red star in panel (a), thus showing good convergence to the ground state in thermodynamic limit. Panel (b) shows the zoom-in plot in the critical regime. The difference between different sizes only appears in the third decimal place. $\mathcal{E}^{b}$ peaks near the QCP slightly towards the DSM phase at around $V=1.2$.
II.2 $\mathcal{E}^{f}$ in finite temperature
Figure S9: Nearest neighbor entanglement negativity $\mathcal{E}^{f}$ in finite temperature The changes of $\mathcal{E}^{f}$ against (a) temperature $T$ and inverse temperature $\beta$. Dashed lines are fitted power-law decay curves at corresponding $V$. The two panels share the same legend in (a).
Fig. S9 shows the nearest neighbor LN $\mathcal{E}^{f}$ at finite temperature, with $V=0.20$ (DSM phase), $V=1.28$ (near QCP), and $V=3.00$ (CDW phase) respectively. The result of high-temperature fitting yields powers close to 2, which is the theoretical prediction in free case by Ref [32]. a
Figure S10: Separated entanglement negativity $\mathcal{E}^{f}$ in finite temperature The changes of $\mathcal{E}^{f}$ against inverse temperature $\beta$, with $A_{1}$ and $A_{2}$ separated with bond-distance $r$ being (a) 3 and (b) 5. Dashed lines are fitted power-law decay curves at corresponding $V$.
If one further separate the two subregions, the power will increase as $2r$ as shown in the fitting in Fig. S10. In high temperature, the fermionic Green’s function expands as
$G=1-\frac{1}{1+e^{-\beta H}}=\frac{1}{2}-\frac{\beta}{4}H+\frac{\beta^{3}}{48}H^{3}-\frac{\beta^{5}}{480}H^{5}+\ldots,$
and the lowest order attained by any $\langle c^{\dagger}_{i}c_{j}\rangle$ depends on the minimum number of hopping terms connecting them, which is simply the bond-distance $r_{ij}$. That is, in cases like ours where only nearest hopping exists, one has
$\langle c^{\dagger}_{i}c_{j}\rangle\sim\beta^{r_{ij}}=1/T^{r_{ij}}$
In high temperature, $\mathcal{E}^{f}$ goes as $1/G_{ij}^{2}\sim 1/T^{2r_{ij}}$ as confirmed by fitting in Fig. S9 and S10
Figure S11: Entanglement negativity and finite-size scaling in different temperatures and system sizes. (a) $\mathcal{E}^{b}$ with different system sizes in $\beta=0.75,1,2,6$. The stars in different colors are the free fermion ED results done in $L=1000$ for corresponding temperatures. The triangles indicate the maxima of $\mathcal{E}^{f}$ along the $V$ axis (the purple dots in Fig. 2 in the main text).(b) finite-size scaling of the order parameter $M^{2}$ against V. Crosses in different colors indicate crossing of the re-scaled observable in respective temperatures. The two panels share the same legend.
To further ensure the $L=6$ finite-temperature data used to construct the heatmap in main text has converged to the thermodynamic limit, and to find the locations of finite-temperature phase transition $V_{c}$, we also simulate the system in larger system sizes in a few fixed temperature ($\beta=0.5,0.75,1,2$).
As shown in Fig. S11 (a), the finite-temperature $\mathcal{E}^{f}$ in these sizes has converged well, indicated by the overlapping of curves from different system sizes. For all temperatures, the curves extrapolate well to the free fermion (${V=0}$) ED result.
Fig. S11 (b) shows the result of finite-size scaling with order parameter and the 2d Ising critical exponent $\eta_{\text{Ising}}=\frac{1}{4}$. The phase transition point, indicated by crosses in respective colors, makes up the finite-temperature phase boundary in the heatmap in main text.